# MnBr2 on the graphene on Ir(110) substrate: growth, structure, and super-moiré

### v1

@@ Abstract
Single-layer MnBr2 is grown on graphene (Gr) supported by Ir(110) and investigated using low-energy electron diffraction, scanning tunneling microscopy, and spectroscopy. The structure and epitaxial relationship with the substrate are systematically characterized. The growth morphology strongly depends on the growth temperature, evolving from fractal to dendritic and eventually to compact dendritic–skeletal islands, reflecting changes in the underlying surface diffusion processes. MnBr2 on Gr/Ir(110) constitutes a three-lattice system, giving rise to a super-moiré pattern –a moiré of moirés. Due to the involvement of lattices of differing symmetries and the partial electronic transparency of Gr, a "virtual" moiré formed by MnBr2 and Ir(110) contributes to the super-moiré formation. Ab initio calculations play a crucial role in understanding the complexity of super-moiré. Moreover, the pronounced variation in the apparent height with tunneling conditions for the magnetic insulator is explained based on the measured electronic structure.
@@ I Introduction
Transition metal dihalides (TMDHs) are a class of materials that often exhibit magnetic ordering in their single-layer form owing to their unique electronic structures and strong spin-orbit coupling [1, 2, 3, 4, 5, 6]. Since the experimental discovery of intrinsic magnetic ordering in two-dimensional materials, single-layer TMDHs have attracted substantial research interest due to their potential as low-dimensional magnets and their applicability in spintronics and quantum technologies [7, 8, 9].
Beyond their intriguing magnetic properties, epitaxially grown single-layer TMDHs on graphene (Gr) undergo spontaneous self-trapping of charges, forming quasiparticles with extremely long lifetimes, identified as polarons [10, 11, 12, 13]. The formation of these quasiparticles is of fundamental and technological significance because they can mediate not only charge but also spin information in magnetic materials. However, the exact nature of these quasiparticles remains debatable because of their exceptionally long lifetimes. Some studies attribute them to small polarons [10, 11, 13, 14], whereas others explain them as solitons [15, 16]. Because these quasiparticles are observed in epitaxially grown single-layers TMDHs, understanding their exact nature requires a detailed analysis of the TMDH-Gr epitaxial relationship.
@@
The investigation of single-layer TMDHs is often limited to first-principles calculations, with few experimental validations. An example lacking experimental realization is 2D manganese dibromide (MnBr2). For bulk MnBr2 a double-row striped antiferromagnetic ordering in the layers below 2.15 K was measured. This order becomes sinusoidally modulated between 2.15 K and 2.35 K before turning paramagnetic [17, 18, 19, 20]. Bulk MnBr2 crystallizes in a layered trigonal CdI2-type structure (space group: $\rm{P\bar{3}m1}$), where the layers are separated by van der Waals (vdW) gaps. In each layer, Mn2+ ions occupy an octahedral site coordinated by six Br- ions in an edge-sharing arrangement. Above 623 K bulk MnBr2 undergoes a phase transition in the layer stacking sequence, transforming the trigonal CdI2-type to rhombohedral CdCl2-type structure (space group: $\rm{R\bar{3}m}$) [21]. Due to the Mn2+ high-spin d5 configuration in dihalides, Mn possesses the highest magnetic moment of $\approx 5\mu_{\mathrm{B}}$ among the TMDHs. The combination of the low electronegativity of Mn (enhancing ionic interactions with halides) and its half-filled 3d5 configuration (suppressing electron hopping) creates favorable conditions for charge localization and enhances electron-phonon coupling. These characteristics suggest that polarons could also form within Mn dihalides.
[…]
@@
The preparation of single-layer MnBr2 has not yet been realized. Preparation by exfoliation may be challenging because the strength of the in-plane bonds of TMDHs is weaker than that of materials such as Gr or transition metal dichalcogenides. Presumably for this reason, most experimental studies on single layers of TMDHs have been conducted using samples grown by molecular beam epitaxy (MBE) [26, 27, 28, 29, 30, 11, 10, 31, 32, 33, 12, 34, 35, 36]. Therefore, MBE of MnBr2 is also the method of choice used in this study.
@@
In recent years, high-order moiré patterns that emerge from the geometrical interference of more than two lattices, so-called super-moiré or moiré of moiré, have captured significant research interest within the condensed matter physics community because they provide an exceptional platform for realizing exotic quantum states [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. For instance, super-moiré systems can host topological flat bands [49, 50, 51, 47, 48] that enhance the electron-electron interactions, leading to the emergence of strongly correlated states, such as unconventional superconductors [52], fractional Chern insulators [53], quantum spin liquids [54], and Wigner crystals [55]. In $k$-space, a moiré results from an integer linear combination of vectors $\vec{k}[\mathrm{A}]$ and $\vec{k}[\mathrm{B}]$ belonging to two different lattices A and B, resulting in moiré vectors $\vec{k}[\mathrm{A-B}]$. Consequently, a super-moiré is considered an integer linear combination of two different sets of moiré vectors that originate from two different moirés [41, 42]. This requires a minimal set of three lattices, A, B, and C, with sets $\vec{k}[\mathrm{A-B}]$ and $\vec{k}[\mathrm{B-C}]$. To date, super-moiré systems have been mostly limited to twisted triple-layer Gr [45, 46, 47, 40] or heterostructures between Gr and hexagonal boron nitride [41, 42, 43, 44, 38], where all the lattices have hexagonal symmetries.
Here, we investigate the case of MnBr2/Gr/Ir(110), in which two lattices are hexagonal (Gr and MnBr2), but one is rectangular, namely the surface lattice of Ir(110). The resulting super-moiré not only exhibits intriguing complexity, but is also formed differently than outlined above. Instead of resulting from the combination of moiré vectors $\vec{k}[\mathrm{A-B}]$ and $\vec{k}[\mathrm{B-C}]$, its dominating Fourier coefficient results from a vector $\vec{k}[\mathrm{A-C}]$, that is, in the present case $\vec{k}[\mathrm{MnBr_{2}/Ir(110)}]$. This is surprising, since Ir(110) and the MnBr2 layer are not in contact with each other. Nevertheless, this study provides a tractable explanation for this phenomenon.
@@ II.1 Structure
Figure 1: Structural characterization of MnBr2. (a) Contrast-inverted 90 eV LEED pattern of MnBr2 after the growth of 1.1 ML at 400 K on Gr/Ir(110). First-order Ir, Gr, and MnBr2 reflections are encircled in magenta, green, and orange, respectively. Reciprocal Ir, Gr and MnBr2 primitive translations are indicated. (b) STM topography corresponding to the sample in (a). (c) Atomically resolved STM image of MnBr2. (c) and (d) Top and side view ball models of the structure and bilayer stackings of MnBr2. All subfigures are oriented as indicated in (a), with the [001]-direction of Ir being horizontal. STM images are obtained at 300 K with (b) $V_{\mathrm{b}}$ = -2 V, 50 pA and (c) $V_{\mathrm{b}}$ = -2 V, 500 pA. Image sizes: (b) 180 nm $\times$ 200 nm, and (c) 7.5 nm $\times$ 7.5 nm.
@@
The first insight into the MnBr2 structure and epitaxial relationship with Gr/Ir(110) is provided through the low-energy electron diffraction (LEED) pattern in Figure 1(a), taken after the growth of 1.1 ML at 400 K. The LEED pattern displays the first-order diffraction spots of the rectangular lattice of unreconstructed Ir(110) [encircled in magenta, with primitive reciprocal translations $\vec{a}^{*}_{Ir,1}$ and $\vec{a}^{*}_{Ir,2}$], hexagonal Gr lattice [encircled in green, with primitive reciprocal translations $\vec{a}^{*}_{Gr,1}$ and $\vec{a}^{*}_{Gr,2}$], and hexagonal MnBr2 lattice [encircled in orange, primitive reciprocal translations $\vec{a}^{*}_{MnBr_{2},1}$ and $\vec{a}^{*}_{MnBr_{2},2}$)]. Gr and Ir(110) are epitaxially aligned with a Gr zigzag direction parallel to $001]_{\mathrm{Ir}} [$56]. MnBr2 forms a hexagonal 2D lattice epitaxially aligned to Gr, with the two hexagonal lattices rotated by 30 with respect to each other. However, the epitaxy is not perfect, as manifested by the azimuthal smearing of the MnBr2 diffraction spots and the presence of a faint, barely visible diffraction ring (highlighted by the dotted circle segment and black arrow). Using the Ir diffraction spots as a reference, the analysis of multiple LEED patterns yields a MnBr2 lattice parameter of 3.90 Å $\pm$ 0.01 Å. This value is in excellent agreement with both the experimental bulk MnBr2 lattice constant of 3.873 Å [57] and the theoretically predicted values for single-layer MnBr2 [3, 25, 24]. Other spots not mentioned and visible in the LEED pattern can be traced back to the moiré of Gr on Ir(110) [56].
@@
Figure 1(b) presents a scanning tunneling microscopy (STM) overview topograph of the same sample characterized by LEED, revealing two large first-layer MnBr2 islands partially overlaid by smaller coalesced second-layer islands on top. The islands exhibit a triangular morphology with mostly straight edges, consistent with the C3v symmetry expected for a single layer of MnBr2. At $V_{\mathrm{b}}=-2.0$ V, the apparent height of the first-layer island is measured to be 4.15 Å [see Figure S1 of the supporting information (SI)]. This value is lower than the expected apparent height of the TMDH on Gr. The underlying reason for this discrepancy is discussed in detail in the Apparent Height and Density of States section. The contrast-enhanced insets of the first-layer islands in Figure 1(b) also display different stripe patterns. The stripe patterns are the super-moirés of single-layer MnBr2 with the underlying substrate, which are discussed in more detail in the Super-moiré section. The different orientations and appearances of the two stripe patterns result from the different orientations of the MnBr2 islands. They are rotated by $\approx 6^{\circ}$ with respect to each other and thus have different orientations with respect to the substrate.
@@
Figure 1(c) shows an atomically resolved STM image of single-layer MnBr2, which confirms a hexagonal lattice that is consistent with the LEED results. In addition to the atomic-scale corrugation of $\approx$ 0.2 Å, the image exhibits nanoscale modulations manifested as narrow and wide vertical stripes. The vertical stripes are the same as those visible in the contrast-enhanced insets of Figure 1(b).
The second-layer islands in Figure 1(b) have two orientations with respect to the first-layer islands. Either their triangular envelope aligns with their base or opposes it. From the analysis of many STM topographs, it is found that aligned and opposing envelopes of second-layer islands are present in nearly equal proportions. As visualized in the ball model of Figure 1(d), the aligned envelopes correspond to bulk MnBr2 which crystallizes in the CdI2 structure [17], also denoted 1T-MnBr2. In 1T-MnBr2, layers composed of octahedrally coordinated Mn ion planes sandwiched between Br ion planes are stacked exactly on top of each other. The envelopes of opposite orientations indicate a stacking fault. Deviations from 1T stacking are well known for bulk MnBr2: above 623 K MnBr2 in the CdCl2 structure is the stable phase, that is, the layers are rhombohedrally stacked, giving rise to a tripled c-axis parameter [21]. This implies small free energy differences for variations in stacking, consistent with the formation of stacking faults under the non-equilibrium conditions of MBE growth. In agreement with our observations, DFT calculations in the BiDB database [58] found identical energies for the two experimental MnBr2 stackings within the precision of DFT, that is, an energy difference of only 0.1 meV /Å2.
Figure 2: (a)-(c) STM topographs of MnBr2 on Gr/Ir(110) after deposition of $\approx 0.6$ ML at 80 K, 300 K, and 400 K, respectively. The sample in (a) was briefly annealed to $\approx 150$ K during transfer to the STM and then imaged at 77 K. Images in (b) and (c) were acquired at 300 K. The white arrows in (b) indicate the dendritic growth directions of the primary, secondary, and ternary dendrite branches. STM imaging parameters: $V_{\mathrm{b}}=-2$ V and $I_{\mathrm{t}}=50$ pA. Image sizes: 160 nm $\times$ 160 nm.

### v2

@@ Abstract
Single-layer MnBr2 is grown on graphene (Gr) supported by Ir(110) and investigated using low-energy electron diffraction, scanning tunneling microscopy, and spectroscopy. The structure and epitaxial relationship with the substrate are systematically characterized. The growth morphology strongly depends on the growth temperature, evolving from fractal to dendritic and eventually to compact dendritic–skeletal islands, reflecting changes in the underlying surface diffusion processes. The pronounced variation in the apparent height with tunneling conditions for the magnetic insulator is explained based on the measured electronic density of states. MnBr2 on Gr/Ir(110) constitutes a three-lattice system, giving rise to a super-moiré pattern – a moiré of moirés. The super-moiré of MnBr2/Gr/Ir(110) is unique, as it involves a virtual moiré of MnBr2 with the Ir(110) surface lattice – two lattices not in contact with each other. Using a careful Fourier analysis, the known properties of Gr/Ir(110), and the results of ab initio calculations, the origin of the virtual moiré is uncovered and related to the inhomogeneous binding of Gr to Ir(110). Comparative experiments with MnBr2 on Gr/Ir(111) show similar growth and structure, but highlight the unique properties of the MnBr2/Gr/Ir(110) super-moiré.
Super-moiré, Moiré of moiré, Transition metal dihalides, MnBr2, 2D magnetic material.
@@ I Introduction
Transition metal dihalides (TMDHs) are a class of materials that often exhibit magnetic ordering in their single-layer form owing to their unique electronic structures and strong spin-orbit coupling [Botana2019, An2020, Bo2024, Kulish2017, McGuire2017, Basak2023]. Since the experimental discovery of intrinsic magnetic ordering in two-dimensional materials, single-layer TMDHs have attracted substantial research interest due to their potential as low-dimensional magnets and their applicability in spintronics and quantum technologies [Huang2017, Elahi2022, Wang2022]. Beyond their intriguing magnetic properties, epitaxially grown single-layer TMDHs on graphene (Gr) undergo spontaneous self-trapping of charges, forming quasiparticles with extremely long lifetimes, identified as polarons [cai2023, liu2023, Miao2025, Liang2025, Yao2024, Hao24, Song25]. Because these quasiparticles are observed in epitaxially grown single-layer TMDHs, understanding their exact nature requires a detailed analysis of the TMDH-Gr epitaxial relationship.
@@
The investigation of single-layer TMDHs is often limited to first-principles calculations, with few experimental validations. An example lacking experimental realization is 2D manganese dibromide (MnBr2). Bulk MnBr2 crystallizes in a layered trigonal CdI2-type structure (space group: $\rm{P\bar{3}m1}$), where the layers are separated by van der Waals (vdW) gaps. In each layer, Mn2+ ions occupy an octahedral site coordinated by six Br- ions in an edge-sharing arrangement. At low temperatures, for bulk MnBr2 a double-row striped antiferromagnetic ordering in the layers was measured [Wollan58, farge1976, Iio1990, Sato1994].
Single-layer MnBr2 has been investigated theoretically using density functional theory (DFT) [Kulish17, Botana19, Luo20, Bo2024, Jesus25]. The different calculations agree that single layers are stable in the trigonal phase, similar to the structure of a single layer in bulk crystals. According to calculations, antiferromagnetic order is also present in the single-layer. The predicted band gap of the magnetic insulator ranges from 2.61 eV [Bo2024] to 4.21 eV [Luo20].
@@
The preparation of single-layer MnBr2 has not yet been realized. Preparation by exfoliation may be challenging because the strength of the in-plane bonds of TMDHs is weaker than that of materials such as Gr or transition metal dichalcogenides. Presumably for this reason, most experimental studies on single layers of TMDHs have been conducted using samples grown by molecular beam epitaxy (MBE) [Zhou2020, Cai2020, Aguirre2024, Xiang2024, Hadjadj2023, liu2023, cai2023, Peng2020, Amini2024, wang2024, miao2023, Bikaljević2021, Terakawa2023, Cai2021]. Therefore, MBE of MnBr2 is also the method of choice used in this study.
In recent years, high-order moiré patterns that emerge from the interference of more than two lattices, so-called super-moiré or moiré of moiré patterns, have captured significant research interest within the condensed matter physics community because they provide an exceptional platform for realizing exotic quantum states [Zhang2021, Ma2025, Kwan2025, Xie2025, Wang19, Andelkovic20, Hesp24, Lai2025, Craig24, Park25, Xia2025, Guerci2024]. For instance, super-moiré systems can host topological flat bands [AlEzzi2024, Hao2024_flatband, Makov2024, Xia2025, Guerci2024] that enhance the electron-electron interactions, leading to the emergence of strongly correlated states, such as unconventional superconductors [Balents2020], fractional Chern insulators [Liu2024], quantum spin liquids [Guangze2021], and Wigner crystals [Shayegan2022]. In $k$-space, a moiré results from an integer linear combination of vectors $\vec{k}[\mathrm{A}]$ and $\vec{k}[\mathrm{B}]$ belonging to two different lattices A and B, resulting in moiré vectors $\vec{k}[\mathrm{A/B}]$. Consequently, a super-moiré is considered an integer linear combination of two different sets of moiré vectors that originate from two different moirés [Wang19, Andelkovic20]. This requires a minimal set of three lattices, A, B, and C, with sets $\vec{k}[\mathrm{A/B}]$ and $\vec{k}[\mathrm{B/C}]$. To date, super-moiré systems have been mostly limited to twisted triple-layer Gr [Craig24, Park25, Xia2025, Xie2025] or heterostructures between Gr and hexagonal boron nitride [Wang19, Andelkovic20, Hesp24, Lai2025, Ma2025], where all the lattices have hexagonal symmetries.
@@
The case of MnBr2/Gr/Ir(110) expands the concept and knowledge of super-moirés. Two of the three lattices are hexagonal (Gr and MnBr2), but one is rectangular, namely the surface lattice of Ir(110). The resulting super-moiré not only exhibits intriguing complexity, but is also formed differently than anticipated previously. Instead of resulting from the combination of moiré vectors $\vec{k}[\mathrm{A/B}]$ and $\vec{k}[\mathrm{B/C}]$, its dominating Fourier coefficient results from a vector $\vec{k}[\mathrm{A/C}]$, that is, in the present case $\vec{k}[\mathrm{MnBr_{2}/Ir(110)}]$. This is surprising since Ir(110) and the MnBr2 layer are not in contact with each other. Hence, the effect is coined as virtual moiré. The presence of a virtual moiré in super-moiré formation is also unique, as highlighted by comparative experiments for MnBr2 on Gr/Ir(111), which otherwise show a rather similar growth and structure of MnBr2.
@@ II.1 Structure
Figure 1: Structural characterization of MnBr2. (a) Contrast-inverted 90 eV LEED pattern of MnBr2 after the growth of 1.1 ML at 400 K on Gr/Ir(110). First-order Ir, Gr, and MnBr2 reflections are encircled in magenta, green, and orange, respectively. Reciprocal Ir, Gr, and MnBr2 primitive translations are indicated. (b) STM topography corresponding to the sample in (a). Contrast enhanced insets display striped moirés highlighted by white lines. (c) Atomically resolved STM image of MnBr2. White line highlights moiré stripe, as in (b). (d) Top and (e) side view ball models of the structure and bilayer stackings of MnBr2. All subfigures are oriented as indicated in (a), with the [001]-direction of Ir being horizontal. MnBr2 unit cells are indicated as black rhombuses in (c) and (d). STM images are obtained at 300 K with (b) $V_{\mathrm{b}}$ = -2 V, 50 pA and (c) $V_{\mathrm{b}}$ = -2 V, 500 pA. Image sizes: (b) 180 nm $\times$ 200 nm, and (c) 7.5 nm $\times$ 7.5 nm.
@@
The first insight into the MnBr2 structure and epitaxial relationship with Gr/Ir(110) is provided through the low energy electron diffraction (LEED) pattern in Figure 1(a), taken after the growth of 1.1 ML at 400 K. The LEED pattern displays the first-order diffraction spots of the rectangular lattice of unreconstructed Ir(110) [encircled in magenta, with primitive reciprocal translations $\vec{a}^{*}_{Ir,1}$ and $\vec{a}^{*}_{Ir,2}$], the hexagonal Gr lattice [encircled in green, with primitive reciprocal translations $\vec{a}^{*}_{Gr,1}$ and $\vec{a}^{*}_{Gr,2}$], and the hexagonal MnBr2 lattice [encircled in orange, primitive reciprocal translations $\vec{a}^{*}_{MnBr_{2},1}$ and $\vec{a}^{*}_{MnBr_{2},2}$]. Gr and Ir(110) are epitaxially aligned with a Gr zigzag direction parallel to $001]_{\mathrm{Ir}} [$Kraus22]. MnBr2 forms a hexagonal 2D lattice epitaxially aligned to Gr, with the two hexagonal lattices rotated by 30 with respect to each other. However, the epitaxy is not perfect, as manifested by the azimuthal smearing of the MnBr2 diffraction spots and the presence of a faint, barely visible diffraction ring (highlighted by the segment of the dotted circle and the black arrow). The smearing of the diffraction spots becomes broader and the diffraction ring more intense when the growth temperature is lower. As typical in van der Waals epitaxy, the lower the growth temperature, the larger the orientation scatter. Using the Ir diffraction spots as a reference, analyzing multiple LEED patterns yields a MnBr2 lattice parameter of 3.90 Å $\pm$ 0.01 Å. This value is in excellent agreement with both the experimental bulk MnBr2 lattice constant of 3.873 Å [Ronda87] and the theoretically predicted values for single-layer MnBr2 [Bo2024, Jesus25, Luo20]. Other spots not mentioned and visible in the LEED pattern can be traced back to the moiré of Gr with Ir(110) [Kraus22].
@@
Figure 1(b) presents a scanning tunneling microscopy (STM) overview topograph of the same sample characterized by LEED, revealing two large first-layer MnBr2 islands partially overlaid by smaller coalesced second-layer islands on top. The islands exhibit a triangular morphology with mostly straight edges, consistent with the C3v symmetry expected for a single layer of MnBr2. At $V_{\mathrm{b}}=-2.0$ V, the apparent height of the first-layer island is measured to be 4.15 Å [see Figure S1 of the supporting information (SI)]. This value is lower than the expected geometric height of MnBr2 on Gr, since the bulk interlayer spacing of MnBr2 is 6.272 Å [Wollan58]. The underlying reason for this discrepancy is discussed in detail in the Section Apparent Height and Density of States. The contrast-enhanced insets of the first-layer islands in Figure 1(b) also display different stripe patterns. The stripe patterns are the moirés of single-layer MnBr2 with the underlying substrate, which are discussed in more detail in the Section Super-Moiré. The different orientations and appearances of the two stripe patterns result from the different orientations of the MnBr2 islands consistent with the smeared LEED spots and the faint diffraction ring of MnBr2. The islands are rotated by $\approx 6^{\circ}$ with respect to each other.
@@
Figure 1(c) shows an atomically resolved STM image of single-layer MnBr2, which confirms a hexagonal lattice that is consistent with the LEED results. In addition to the atomic-scale corrugation of $\approx$ 0.2 Å, the image exhibits nanoscale modulations. Most prominent are bright nearly vertical stripes, which are the same as those visible in the contrast-enhanced insets of Figure 1(b). A small number of dark spots of atomic size are present in the atomically resolved image. Since their concentration does not show a dependence on growth or annealing temperature, they could result from a slight Br loss during the molecular sublimation in the Knudsen cell at 670 K and be Br vacancies.
The second-layer islands in Figure 1(b) have two orientations with respect to the first-layer islands. Either their triangular envelope aligns with their base or opposes it. From the analysis of many STM topographs, it is found that aligned and opposing envelopes of second-layer islands are present in nearly equal proportions. As visualized in the ball model of Figure 1(d), the aligned envelopes correspond to bulk MnBr2 which crystallizes in the CdI2 structure [Wollan58], also denoted 1T-MnBr2. In 1T-MnBr2, layers composed of octahedrally coordinated Mn ion planes sandwiched between Br ion planes are stacked exactly on top of each other. The envelopes of opposite orientations indicate a stacking fault. Deviations from 1T stacking are well known for bulk MnBr2: above 623 K MnBr2 in the CdCl2 structure is the stable phase, that is, the layers are rhombohedrally stacked, giving rise to a tripled c-axis parameter [Schneider1992]. This implies small free energy differences for variations in stacking, consistent with the formation of stacking faults under the non-equilibrium conditions of MBE growth. In agreement with our observations, DFT calculations in the BiDB database [Pakdel24] found identical energies for the two experimental MnBr2 stackings within the precision of DFT, that is, an energy difference of only 0.1 meV /Å2.
Figure 2: (a)-(c) STM topographs of MnBr2 on Gr/Ir(110) after deposition of $\approx 0.6$ ML at 80 K, 300 K, and 400 K, respectively. The sample in (a) was briefly annealed to $\approx 150$ K during transfer to the STM and then imaged at 77 K. Inset in (a) uses double gray scale to make the moiré on Gr/Ir(110) and the moiré on the MnBr2 islands simultaneously visible. Images in (b) and (c) were acquired at 300 K. The white arrows in (b) indicate the dendritic growth directions of the primary, secondary, and ternary dendrite branches. STM imaging parameters: (a) $V_{\mathrm{b}}=-1$ V and $I_{\mathrm{t}}=50$ pA, (b,c) $V_{\mathrm{b}}=-2$ V and $I_{\mathrm{t}}=50$ pA Image sizes: 160 nm $\times$ 160 nm.